how can I show my work for the problem x4+7x2=18
Answer by stanbon(75887)

how can I show my work for the problem x4+7x2=18
--------
x^4 + 7x^2 - 18 = 0
------
(x^2+9)(x^2-2) = 0
----
x = +/-3i or x = +-isqrt(2)
================================
Cheers,
Stan H.